Read Chapter 7: Cindy is a student at Valencia planning to enter Nursing School Cindy is 5 feet 2 inches She weighs 158 pounds Her waist measures 42 inches What is her BMI? Using the BMI index, state what is her health risk? Her wait measures 42 inches what does this indicate? Cindy has started to diet by cutting 500 kcal /day. She used to eat 2100 kcal /day. She has started to use the treadmill at a local YMCA. She burns 500 kcal on the tread mill/day Therefore she is losing “x” amount of calories by dieting and “x” amount of calories by exercising. (Pay attention to this sentence) Question: After 1 month how many pounds do you calculate Cindy could lose? Please show your calculation What would be her weight in pounds in one month? Cindy continues to cut 500 Kcal/day (dieting) and burning 500 kcal/day, using treadmill. What would be her weight in pounds after 6 months? Use her initial weight 158 pounds to compute her loss of weight after six month. HINT: 3500kcal must be burned in order to lose 1 pound of fat Weight loss is calculated by pounds lost/ week 1 month= 4 weeks 6 month=24 weeks
Answer #1

Answer :

* Cindy's height is 5 feet 2 inches, so in meters the height is

" 1.57 meters "

* weight is 158 pounds, if we converts into kgs the weight is

" 71.66 kgs "

* the BMI formula is = weight inkgs / height metre square.

* 71.66/ 1.57×1.57

= 71.66/ 2.46

= 29.13 kg per meter square.

* according to her BMI she is having over weight, because 18 to 24.9 is the normal

* she is having health risk of developing the cardiovascular diseases, diabetes mellitus, sleep problems, social and emotional problems.

* her waist circumference is 42 inches it means 106. 68 cm.

* it indicates she is high risk for development of cardiovascular diseases and diabetes mellitus.

* the Cindy loosing 1000 k. Cal per day both diet and exercise.

* that means per week she is loosing 1000× 7 = 7000 k. Cal.

* that means 3500 k. Call burns means she can loose 1 pound weight.

* so here 7000 k. Cal means 2 pounds she is loosing per week.

* so 2 pounds per weeks means per month she can loose 8 pounds, because per month 4 weeks.

* so 2 × 4 = 8 pounds in a month she is loosing.

* so after 6 months she can loos 48 pounds because, = 6 months × 8 pounds loss per month, so 6 × 8 = 48 pounds.

* finally after 6 months her weight is 158 - 48 = 110 pounds.

* now her BMI is 20. 28 kg per meter square ,it is in correct now.

* because 110 pounds means 49.89 kgs,

* BMI = 49.89 / 2.46 = 20.28 kg per meter square.